# Taming Data and Transformers for Audio Generation

### v2

@@ D.2 GenAu
Appendix E Additional Results
@@
E.1 Additional Audio Captioning Evaluation
In Tab. 9 we show qualitative results of the captions produced by our method and compare them with state-of-the-art AAC methods. See the Website for qualitative results accompanied by the original audio. While ENCLAP (Kim et al., 2024b) and CoNeTTE (Étienne Labbé et al., 2023) tend to produce short captions, our method produces the most descriptive captions, capturing the most amount of elements from the ground truth audio, an important capability to allow high-quality audio generation (Shi et al., 2020).
E.2 Additional Audio Generation Evaluation
@@
In Tab. 11, we ablate the 1D-VAE bottleneck size in terms of reconstruction loss and performance of a subsequently trained latent audio diffusion model, in terms of FAD, FD, and IS. Similarly to the phenomenon observed in the image and video generation domain (Gupta et al., 2023; Esser et al., 2024), we observe that a larger number of channels allocated to the latent space results in lower reconstruction losses, but making the latent space more complex, hindering generation quality. We adopt 64 1D-VAE channels for all our experiments.

### v3

@@ E.2 GenAu
Appendix F Discussion with Concurrent work
F.1 Text-conditioned audio generation
Recently, Stable Audio Open (Evans et al., 2024b) introduced a 1.32B-parameter model capable of generating variable-length stereo audio clips at 44.1 kHz. This model leverages a latent diffusion approach with a DiT (Peebles & Xie, 2023) as its diffusion backbone, similar to prior work such as Make-An-Audio 2 (Huang et al., 2023a). In contrast, GenAu employs a FiT architecture. In Tab. 5, we show the superiority of our FiT-based approach over DiT by showing that GenAu-S is consistently preferred over a 937M-parameter DiT-based baseline (Make-An-Audio 2 (Huang et al., 2023a)) when trained on comparable data settings (i.e.without recaptioning) at a smaller scale (493M parameters). Additionally, Stable Audio Open proposes directly encoding audio clips using a variational autoencoder (VAE) with a ResNet-like architecture, which is particularly effective for higher-resolution audio generation. In contrast, our work adopts previous approaches (Huang et al., 2023a; Liu et al., 2023b) and uses a Mel-spectrogram representation due to its simplicity. GenAu, being a latent model, can readily benefit from improved latent audio representations, such as those employed by Stable Audio Open.
F.2 Audio captioning
A concurrent work, SOUND-VECAPS (Yuan et al., 2024), and Auto-ACD (Sun et al., 2024), propose prompting a pretrained large language model with multimodal information. SOUND-VECAPS utilizes visual captions generated by a pretrained visual captioner (Wang et al., 2025) alongside audio captions from a pretrained audio captioner, ENCLAP (Kim et al., 2024b), to produce more complex captions, showing significant improvements in the downstream task of audio generation. This aligns with our approach of incorporating visual captions in the audio captioning task. However, unlike these methods, which rely solely on pretrained models, we integrate visual information directly into the training process of the audio captioner. This enables a more dynamic and context-aware incorporation of visual information in the audio captioning task.
Additionally, there has been a recent trend toward training large audio-language models (Ghosh et al., 2024b; Kong et al., 2024; Gong et al., 2023; Deshmukh et al., 2023) and utilizing them for audio captioning in zero-shot settings. While promising in the pursuit of general-purpose models, their reported results on audio captioning remain inferior to state-of-the-art automatic audio captioning (AAC) methods. Consequently, we opt to train a dedicated AAC model, AutoCap, to achieve the highest-quality captions for our proposed dataset, AutoReCap.
Appendix G Additional Results
@@
G.1 Additional Audio Captioning Evaluation
In Tab. 9 we show qualitative results of the captions produced by our method and compare them with state-of-the-art AAC methods. See the Website for qualitative results accompanied by the original audio. While ENCLAP (Kim et al., 2024b) and CoNeTTE (Labb et al., 2024) tend to produce short captions, our method produces the most descriptive captions, capturing the most amount of elements from the ground truth audio, an important capability to allow high-quality audio generation (Shi et al., 2020).
G.2 Additional Audio Generation Evaluation
@@
In Tab. 11, we ablate the 1D-VAE bottleneck size in terms of reconstruction loss and performance of a subsequently trained latent audio diffusion model, in terms of FAD, FD, and IS. Similarly to the phenomenon observed in the image and video generation domain (Gupta et al., 2024; Esser et al., 2024), we observe that a larger number of channels allocated to the latent space results in lower reconstruction losses, but making the latent space more complex, hindering generation quality. We adopt 64 1D-VAE channels for all our experiments.
G.3 Additional HTSAT Embedding Extraction Evaluation
Table 12: Ablation of token replication factors for the HTSAT embeddings extraction procedure of (Mei et al., 2024a) on the AudioCaps test split. Larger token replication factors consistently improve performance due to the related compute increase in the downstream model.
Tokens Count Replication Factor CIDEr BLEU1 BLEU4 ROUGEL
HTSAT-BART 32 1x 73.7 68.6 25.0 49.7
HTSAT-BART 256 8x 74.4 69.7 26.0 49.8
HTSAT-BART 1024 32x 76.6 71.5 26.3 49.8
AutoCap 32 1x 81.9 71.7 28.9 51.3
AutoCap 1024 32x 82.7 72.5 29.3 52.0
Figure 14: Ablation of token replication factors for the HTSAT embeddings extraction procedure of (Mei et al., 2024a) on the AudioCaps test split for the HTSAT-BART (Mei et al., 2024a) and our AutoCap model.
We perform a series of ablations on HTSAT-BART (Mei et al., 2024a) employing different variants of the procedure of (Mei et al., 2024a) for the extraction of HTSAT embeddings (see Appx. B.1). We consider HTSAT output tokens of shape $32\times 768$ after the averaging operation over the frequency dimension of (Mei et al., 2024a), and apply different token repetition factors to produce embeddings with 32 tokens (no token repetition), 256 tokens (8x token repetition) and 1024 tokens (32x token repetition following (Mei et al., 2024a)). For completeness, we perform the same ablation on our AutoCap, using as input to the Q-Former 32 tokens (no token repetition) and 1024 tokens (32x token repetition). Training hyperparameters of AutoCap are modified to match HTSAT-BART (Mei et al., 2024a) for the purpose of the ablation.
We followed the training procedure of (Mei et al., 2024a) and report evaluation results on the AudioCaps test split for the last obtained checkpoint in Tab. 12 and Fig. 14. As the ablation shows, the token replication operation consistently improves model performance. We attribute this finding to the increased computation in the downstream model caused by it and consequently adopt the best performing 32x token replication embeddings extraction procedure of (Mei et al., 2024a) throughout our work.